HIGH-RESOLUTION SPECTROSCOPY OF HOBr IN THE FAR- AND NEAR-INFRARED

\underline{J. ORPHAL}, Q. KOU, F. KWABIA TCHANA, O. PIRALI, AND J.-M. FLAUD, Laboratoire de Photophysique Moléculaire, Centre National de la Recherche Scientifique, Bât. 350, Centre d’Orsay, 91405 Orsay Cedex, France.

Hypobromous acid, HOBr, is an important species in atmospheric chemistry. In the stratosphere it is mainly produced by the reaction between BrO and HO_{2} but rapidly photolyzed by sunlight. In the marine troposphere it is also formed by heterogeneous reactions on sea-salt aerosols and is thus involved in the transport of bromine from the oceans into the atmosphere. Attempts to measure atmospheric HOBr concentrations have been made^{a} using its far-infrared rotational transitions. In order to provide accurate line positions and intensities in this region we have measured high-resolution Fourier-transform absorption spectra of gaseous HOBr between 100–350 cm^{−1} using the Bruker IFS-120 HR at LPPM in Orsay. We have been able to improve the rotational-vibrational parameters of the ground states^{b} and of the ν_{3} bands of HO^{79}Br and HO^{81}Br around 620 cm^{−1}. For the rotational line intensities we have taken into account the Herman-Wallis effects. Furthermore, we have measured high-resolution absorption spectra of the 2ν_{1} bands around 1.4 μm and determined their rotational-vibrational parameters. These overtone bands are observed to be even stronger than the ν_{3} fundamental bands, in agreement with ab-initio calculations^{c}. The results might be useful for detection of atmospheric HOBr using Cavity Ring-Down Spectroscopy with telecommunication diode-lasers operating at room temperature.

aD. G. Johnson, W. A. Traub, K. V. Chance, and K. W. Jucks, Geophys. Res. Lett. 22, 1373 (1995)

bE. A. Cohen, G. A. McRae, T. L. Tan, R. R. Friedl, J. W. C. Johns, and M. Noël, J. Mol. Spectrosc. 173, 55 (1995)

cK. A. Peterson, J. Chem. Phys. 113, 4598 (2000)